OpenStudy (anonymous):

how do you find zeros of polynomials with a degree greater than 3?

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (anonymous):

Nope we have to use synthetic division then factor

OpenStudy (anonymous):

there is a long formula for degree 3, and even longer one for degree 4 and it has been proven that there can never be one for degree 5 however, if it was cooked up by your math teacher, try the possible integer or rationals roots

OpenStudy (anonymous):

How do you factor it if there is 4 terms(2X cubed-13x squared +10x +25

OpenStudy (anonymous):

best way is to find the zeros first, factor second that is, find one zero first

OpenStudy (anonymous):

\[2x^3-13x^2+10x+25=0\] if it has any rational zeros, they are \(\pm1,\pm5,\pm25\) or \(\pm\frac{1}{2},\pm\frac{5}{2},\pm\frac{25}{2}\)

OpenStudy (anonymous):

try the smaller ones first

OpenStudy (anonymous):

in fact, the easiest one other than 1 (which doesn't work) to try is \(-1\) and it will work then factor as \[(x+1)(\text{whatever})\]